Given f(x)=−x^4+2x^3+10x, determine any intervals where the function is concave down.
Answer #1

Ans tin) = – 44 +243 +lon t(n) = -443+620²710 f(n) = – 1242 + 124 12 (ht - 12 21 (26) flen)- at totottt 1 Conan ten Concare